If p = – 2, find the value of:
- 3p + 4p + 7 …...........

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

From the question it is given that p = -2
Then, substitute the value of p in the question
= (-3 × (-2)2) + (4 × (-2)) + 7
= (-3 × 4) + (-8) + 7
= -12–8 + 7
= -20 + 7
= -13

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Given : p = -2
Substituting value of p in the equation,
-3p²+4p+7
= (-3×4) + (4×-2) + 7
= -12-8+7
= -12-1
= -13
 
Hope it helps !
Thank you ✌️